OpenStudy (anonymous):

Use De Moivre's Theorem to prove the identity:

OpenStudy (anonymous):

\[\cos 3\alpha=4\cos ^{3}\alpha-3\cos \alpha\]

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

simply De Moivre's Theorem :- \((\cos x+ i \sin x)^n=\cos (nx ) +i \sin(nx) \)